# Atmospheric stability sets maximum moist heat and convection in the midlatitudes

### v3

@@ Abstract
This study proposed a theoretical prediction for maximum potential moist heat (MSEs) and convection (CAPE) over midlatitude land, both tightly constrained by a preexisting energy (MSE∗) inversion within the lower free troposphere. The maximum MSE∗ at the energy inversion layer (i.e., MSE) acts as a substantial energy barrier that suppresses the free convection of near-surface air parcels, thereby allowing the warming or moistening of near-surface air and the accumulation of convective instability. This process continues until MSEs reaches MSE, such that the near-surface air parcels are sufficiently energetic to overcome the energy barrier, initiate deep free convection, and terminate surface heating. This theory does not address the mechanisms by which moist heat or convective instability increases, which may be driven by various physical processes (e.g., warm advection, adiabatic warming in subsiding air, or diabatic heating from surface fluxes and radiation) associated with large-scale dynamics such as storm tracks or atmospheric blockings (72, 73, 74, 56, 75, 76, 30, 77, 78, 79, 60, 80). Instead, it focuses on quantifying the upper limit of these increases before free convection is triggered, assuming sufficient heating sources.
A recent study examined a similar topic over midlatitude continents, but focusing on extreme dry heat under the assumption of moist neutrality (45), and suggested that 500-hPa MSE acts as an upper bound for MSEs. However, we have shown that this assumption fails for extreme moist heat over midlatitude land. In fact, the neutrality assumption also breaks down for extreme dry heat in the midlatitudes, as MSEs exceeds MSE in most cases of annual maximum surface air temperature (Fig. S5). Our CAPE scaling framework (Eq.1) reveals that MSE generally approximates the minimum energy of the free troposphere (detailed in ), and the surplus of MSEs over MSE effectively defines the accumulated convective instability beyond moist neutrality. When applied to extreme dry heat conditions, our theory also predicts well the maximum potential intensities of MSEs and CAPEc during extreme dry heat periods (Fig. S5–), thereby generalizing our theory and highlighting the critical role of preexisting low-level energy inversions in setting near-surface maximum heat.
Our study advances previous work by carefully addressing the onset of convection, demonstrating how the atmosphere evolves beyond the neutral condition – where parcel buoyancy and CAPE are zero by definition – in alignment with the stored-energy nature of severe convection in the midlatitudes (46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56). Future work could incorporate other processes (e.g., wind shear (81, 82, 83), entrainments (42, 84, 17), and external forcings (85, 86, 83, 87, 88, 89, 71, 69, 70)) to more precisely determine the onset of convection and, consequently, the termination of heatwaves. For example, a recent study highlighted the critical role of dry air entrainment in suppressing deep convection and intensifying moist heat in tropical and subtropical regions (17), though the entrainment influence may vary strongly with the severity of convection in midlatitude due to the dependence on vertical wind shear (90, 91).
The preexisting MSE∗ inversion within the lower free troposphere is typically marked by a pronounced temperature inversion or a layer where the temperature decreases most gradually with height (61, 62, 63, 15, 64, 65, 55, 56, 66). While this study proposed the theoretical framework based on evolution of local convective instability, the formation of preexisting inversion layers might be non-local due to advection of a warm layer from an adjacent region or the adiabatically warmed air from the subsidence of large-scale circulation (e.g., anticyclones) (92, 56, 61, 65, 15). Thus, it might be worthwhile to investigate the origins and evolutions of the inversion layers from the Lagrangian perspective (76, 93, 94, 55, 74, 56) to provide insight into the formation and potential magnitude of MSE, and thus improve the predictability of local potential moist heat and convection. This is also critical for understanding the role of preexisting lower free tropospheric inversion layers in future climates, as shifts in land cover or regionally uneven warming could affect the formation and transport pathways of inversion layers between regions (95, 96, 97, 98, 66, 53, 56), thereby potentially shaping future spatial patterns of extreme heat and convection (99, 100, 101, 26).
@@
We use 3-hourly near-surface and pressure-level ERA5 reanalysis data with a horizontal grid spacing of 0.5∘ latitude by 0.5∘ longitude during 1980–2022 (36, 37), for land grid points between 35∘N and 75∘N at elevations lower than 1000 m (20854 grid points per year in total). ERA5 reanalysis data are widely used in studies of global temperature extremes and convective environments, as they accurately capture observed temperature and moisture profiles in general (102, 103, 56), including inversion features within the lower free troposphere (64, 66). In this study, calculations for all variables are initially performed at the 3-hourly interval and then sampled based on the daily maximum MSEs to create a dataset of daily maximum moist heat. Our analyses focus on the annual maximum moist heat, obtained by further sampling the dataset given the annual maximum MSEs at each grid point for each year. This process yields 43 samples per land grid point, corresponding to the maximum MSEs for each year during 1980–2022. For analyses related to extreme near-surface dry heat, the data are generated in the same manner but are conditioned on the maximum near-surface dry heat, defined as the maximum 2-m air temperature (Ts).
@@
MSE is defined as the sum of sensible heat (), latent heat (), and geopotential energy (), i.e., , where J kg-1 K-1 is the the specific heat capacity of air at constant pressure, J kg-1 is the latent heat of vaporization, m s-2 is gravitational acceleration, is air temperature, is specific humidity, and is height above sea level (47). Near-surface (denoted by subscript “”) MSE, , is calculated using quantities at 2-m above ground surface, where the geopotential height equals to 2 plus the surface elevation in meters. Saturated (denoted by superscript “”) MSE, , is calculated by replacing with the saturated specific humidity (), which is a function of air temperature and pressure. In an adiabatic process, MSE is generally considered a conserved quantity given hydrostatic balance (47, 104, 105), interconvertible with equivalent potential temperature (106, 107, 35). As a part of MSE, dry static energy (DSE) is the sum of and , which is conserved under dry adiabatic conditions and is thermodynamically equivalent to potential temperature (47, 108). is identified as the maximum MSE∗ within the lower free troposphere, specifically at or above but below 300 hPa.
is calculated based on the conversation of (109, 16). Specifically, for each , we solve the equation for the corresponding , where is the saturated specific humidity at temperature and surface pressure . As is proportional with based on the Clausius–Clapeyron scaling (110), exhibits a nearly one-to-one relationship with (Fig. S1) (16, 15). There is a small variation in WBTs given a value of MSEs, due to dependence on the elevation variance (0–1000 m; Fig. S1). Similarly, we calculate the maximum potential (Fig. 4) by solving from the equation at the time of annual maximum .
@@
Considering an undiluted air parcel adiabatically lifted upward from the near-surface (2 m above ground surface; denoted by subscript “” in derivations) and neglecting the relatively small contribution to density changes due to pressure perturbation and virtual temperature correction(111, 47), the parcel buoyancy at a given height () is proportional to differences in sensible heat between the air parcel and the environment (), where is temperature perturbations of the air parcel (, with subscript “” referring to the air parcel) with respect to the environment () and J kg-1 K-1 is the the specific heat capacity of air at constant pressure. During an adiabatic process, it is common to assume that the air parcel conserves its moist static energy such that . Therefore, the vertical profile of or is given by
@@
For , the air parcel is saturated, such that the parcel specific humidity equals its saturated specific humidity (, with superscript “” referring to a quantity at saturation), which is a function of the parcel temperature and air pressure. As pressure difference between the air parcel and environment is negligible (111, 47), the difference in saturated specific humidity between the air parcel and environment is caused by difference in temperature, and thus can be written as a linear relation to the environmental saturated specific humidity () via . We further assume constant and approximate it by using Clausius-Clapeyron equation, where J kg-1 K-1 is gas constant for water vapor. Hence,
@@
Previous studies have explored the relationship between CAPE and environmental MSE, either under the radiative–convective equilibrium framework (114, 115, 52) or through linear regression methods (116). In contrast, our derivation begins with the fundamental definitions of parcel buoyancy and CAPE, similar to (117) but incorporating minimal assumptions about the geometric characteristics of CAPE, to establish a robust linear relationship between CAPE and key atmospheric parameters near the surface and in the free troposphere. The coefficient in Eq.12 may vary with , , and , which can be precisely adjusted using the full scaling equation (Eq.11) when applied to different regions or datasets. Here, the derived coefficient of 0.22 for high-CAPE environments over midlatitude land aligns closely with the linear regression results (0.22–0.29) reported for high CAPE over continental North America (116).
@@
The 3-hourly ERA5 reanalysis data (36, 37) from 1980–2022 are publicly available at https://rda.ucar.edu/datasets/ds633.0/.

### v4

@@ Abstract
This study proposed a theoretical prediction for maximum potential moist heat (MSEs) and convection (CAPE) over midlatitude land, both tightly constrained by a preexisting energy (MSE∗) inversion in the lower free troposphere. The maximum MSE∗ at the inversion layer (MSE) acts as a substantial energy barrier that suppresses the free convection of near-surface air parcels, thereby allowing near-surface warming or moistening and the accumulation of convective instability. This process can continue until MSEs reaches MSE, such that the near-surface air parcels are sufficiently energetic to overcome the energy barrier, initiate deep free convection, and terminate surface heating. This strong buildup of CAPE, which enables excessive increases in surface moist heat, reflects the stored-energy signature of severe convection in the midlatitudes (49, 59, 50, 51, 52, 53, 54, 55, 56, 57, 58).
Our theory advances previous work on convective constraints of extreme heat by explicitly incorporating CAPE buildup such that the atmosphere can evolve beyond the neutral condition (i.e., parcel buoyancy and CAPE are zero by definition), and thus extends many previous studies commonly assuming moist neutrality at the peak time (6, 19, 17, 48). The buildup of CAPE is a non-negligible factor in predicting surface heat extremes: accumulation of CAPE from 0 to 1000 J kg-1 corresponds an increase of 4.5 K in surface temperatures ( from Eq. 1), assuming other conditions remain constant. One prior study proposed that surface moist heat is limited by the tropopause MSE∗ (6), which aligns with our prediction when the low-level inversion barrier is as strong as the tropopause MSE∗. However, not all convection reaches the tropopause in the midlatitudes, and the tropopause constraint tends to overestimate the maximum intensity of surface moist heat if the low-level inversion energy is weaker than the tropopause MSE∗. In addition, a recent study examined extreme dry heat over midlatitude continents under the assumption of moist neutrality (48), suggesting that 500-hPa MSE acts as an upper bound for MSEs. However, we have shown that the neutrality assumption fails for extreme moist heat over midlatitude land (Fig. 1). In fact, the neutrality assumption also breaks down for extreme dry heat in the midlatitudes, as MSEs exceeds MSE in most cases of annual maximum surface air temperature (Fig. S6). Our CAPE scaling framework (Eq.1) reveals that MSE generally approximates the minimum energy of the free troposphere (detailed in ), and the surplus of MSEs over MSE effectively defines the accumulated convective instability beyond moist neutrality. The dry limit theory (48) further assumes completely dry surface conditions that limit the accumulation of moist convective instability and result in an unrealistically high upper bound on surface temperature constrained by a deep dry adiabatic layer. However, this drying assumption contradicts the hypothesis that convective precipitation terminates heatwaves, as moisture accumulation is essential for the initiation and development of moist convection. In contrast, our theory does not require assumptions about surface humidity and explicitly allows the buildup of moist convective energy, therefore offering a tighter and more physically grounded constraint on surface heat. When applied to extreme dry heat, our theory also predicts well the maximum potential intensities of MSEs and CAPEc (Fig. S6–), which further generalizes the theory.
Within the context of convective constraints on heat extremes, accurately identifying the onset of convection is crucial for determining both the termination and peak intensity of surface heat. Future work could incorporate other processes (e.g., wind shear (77, 78, 79), entrainments (45, 80, 20), and external forcings (81, 82, 79, 83, 84, 85, 76, 74, 75)) to more precisely predict convection onset and, consequently, the termination of heatwaves. For example, a recent study emphasized the role of dry air entrainment in suppressing deep convection and intensifying moist heat in tropical and subtropical regions (20), which is quantified by the production of entraining rate () and 850-hPa saturation deficit (MSE850-MSE). Incorporating sub-cloud entrainment into our framework could further modify the convective constraint on surface moist heat as approximately MSE+(MSE-MSE850). While this may increases the weight of low-level MSE∗, the entraining factor needs careful estimation in the midlatitudes, where vertical wind shear associated with convection is typically stronger than in the tropics (86). In midlatitude severe convection, the strong deep-layer vertical wind shear tends to broaden convection and prevent the influence of entrainment (i.e., smaller under larger deep-layer wind shear) (87, 88, 89).
Implication and outlook
Our work centered on parcel theory to quantify the upper bound on local near-surface moist heat and convection before local free convection is triggered. In most midlatitude land regions, extreme moist heat and convection occur concurrently, and MSE provides a tight and attainable constraint on their maximum intensities. In certain regions (e.g., east of the Caspian Sea, south of the Mediterranean, and along the west coast of the United States), this energetic inversion barrier (MSE) remains largely unbroken, and convective precipitation appears incapable to terminate extreme heat events. In these areas, moist heat and convection extremes co-occur less frequently, and processes other than convective precipitation may terminate extreme heat events, which deserves further future investigation. While our theory does not address the specific mechanisms by which surface moist heat or convective instability increases, previous studies have examined a range of contributing processes, including warm air or moisture advection, adiabatic warming in subsiding air, and diabatic heating from surface fluxes and radiation, associated with large-scale dynamics such as storm tracks and atmospheric blocking (90, 91, 92, 58, 93, 94, 33, 95, 96, 97, 65, 98). Bridging these dynamical drivers with thermodynamical constraints is essential for predicting and understanding how, when, and to what extent heatwave and severe convection can approach theoretical limits, and why such limits are not reached in some regions or events.
Our theory emphasizes the critical role of low-level energy inversion layers in shaping extreme moist heat and severe convection events. These inversion features typically form several days before the peak events and vary weakly over time, therefore offering a valuable opportunity for predicting the maximum intensity of moist heat and convection on synoptic timescales. This predictive potential has long been recognized by Severe Local Storm research and operational communities. A notable example is the elevated mixed layer, characterized by a pronounced low-level temperature inversion formed by warm air advection from upstream elevated terrain, which has been identified as a common synoptic-scale feature contributing to intense severe thunderstorm and tornado activity across central North America (66, 99, 69, 70, 71), South America (68), and Europe (100). Recent studies using Lagrangian trajectory methods and analytical models further suggest that such inversion features are widespread and play a key role in many severe convective events globally (54, 57, 55, 58). Our work quantifies the energetic constraint imposed by inversion layers and extends its utility to extreme moist heat. In practice, monitoring the presence of an explicit low-level MSE∗ inversion could serve as a key signal for timing predictions of the potential intensities of surface moist heat and convection extremes. Meanwhile, understanding the evolutions of MSE and MSE is also crucial for refining these predictions over time. Since both heatwaves and severe convective storms are among the most damaging weather extremes, this theory represents a promising step toward improving the prediction and understanding of these events and their potential compound occurrence in the midlatitudes.
Lastly, our theoretical framework provides a valuable lens for predicting and understanding extreme heat and severe convection under future climates. While low-level energy inversion layers serves as local convective barriers, their formations may stem from non-local processes, such as warm-air advection from an adjacent region (e.g., elevated terrain) or adiabatic warming from subsiding air within large-scale circulations (e.g., anticyclones) (101, 58, 66, 70, 18). A Lagrangian perspective can help trace the origins and transport pathways of inversions (94, 102, 103, 57, 92, 58), offering insight into MSE evolution over time. As the climate changes, land cover change and uneven regional warming may alter inversion formation at the source (104, 105), while changes in large-scale driven dynamics (e.g., elevation-dependent warming and storm track shifts) may influence airmass transport pathways (106, 107, 102, 108, 109, 110). These shifts could affect the intensity, height, persistence and spatial extent of inversion layers (71, 55), thereby reshaping future spatial patterns of extreme heat and convective weather (111, 29, 112, 113). For example, elevated temperature inversions have intensified over the central United States (71), likely associated with the amplified warming upstream over the Rocky Mountains (106, 107), which may raise low-level energy barriers and imply that greater near-surface heating will be required to trigger convection. These changes may be responsible for the amplified moist heat stress over central and Midwest United States(29) and the eastward expansion of severe thunderstorms in this region (113, 112). Climate change may also alter external forcings (e.g., fronts, cyclogenesis, moisture convergence) (114, 115), which can modify the timing of convection onset and drive spatial variations in the divergence between realized and potential intensities of heat and convection extremes. Finally, our framework underscores the need to evaluate model skill in simulating low-level inversions and convection initiation, as the associated biases can substantially impact projections of extreme heat, severe convection, and their compound occurrence (116, 117).
@@
We use 3-hourly near-surface and pressure-level ERA5 reanalysis data with a horizontal grid spacing of 0.5∘ latitude by 0.5∘ longitude during 1980–2022 (39, 40), for land grid points between 35∘N and 75∘N at elevations lower than 1000 m (20854 grid points per year in total). ERA5 reanalysis data are widely used in studies of global temperature extremes and convective environments, as they accurately capture observed temperature and moisture profiles in general (118, 119, 58), including inversion features within the lower free troposphere (69, 71). In this study, calculations for all variables are initially performed at the 3-hourly interval and then sampled based on the daily maximum MSEs to create a dataset of daily maximum moist heat. Our analyses focus on the annual maximum moist heat, obtained by further sampling the dataset given the annual maximum MSEs at each grid point for each year. This process yields 43 samples per land grid point, corresponding to the maximum MSEs for each year during 1980–2022. For analyses related to extreme near-surface dry heat, the data are generated in the same manner but are conditioned on the maximum near-surface dry heat, defined as the maximum 2-m air temperature (Ts).
@@
MSE is defined as the sum of sensible heat (), latent heat (), and geopotential energy (), i.e., , where J kg-1 K-1 is the the specific heat capacity of air at constant pressure, J kg-1 is the latent heat of vaporization, m s-2 is gravitational acceleration, is air temperature, is specific humidity, and is height above sea level (59). Near-surface (denoted by subscript “”) MSE, , is calculated using quantities at 2-m above ground surface, where the geopotential height equals to 2 plus the surface elevation in meters. Saturated (denoted by superscript “”) MSE, , is calculated by replacing with the saturated specific humidity (), which is a function of air temperature and pressure. In an adiabatic process, MSE is generally considered a conserved quantity given hydrostatic balance (59, 120, 121), interconvertible with equivalent potential temperature (122, 123, 38). As a part of MSE, dry static energy (DSE) is the sum of and , which is conserved under dry adiabatic conditions and is thermodynamically equivalent to potential temperature (59, 124). is identified as the maximum MSE∗ within the lower free troposphere, specifically at or above but below 300 hPa.
is calculated based on the conservation of (125, 19). Specifically, for each , we solve the equation for the corresponding , where is the saturated specific humidity at temperature and surface pressure . As is proportional with based on the Clausius–Clapeyron scaling (126), exhibits a nearly one-to-one relationship with (Fig. S1) (19, 18). There is a small variation in WBTs given a value of MSEs, due to dependence on the elevation variance (0–1000 m; Fig. S1). Similarly, we calculate the maximum potential (Fig. 4) by solving from the equation at the time of annual maximum .
@@
Considering an undiluted air parcel adiabatically lifted upward from the near-surface (2 m above ground surface; denoted by subscript “” in derivations) and neglecting the relatively small contribution to density changes due to pressure perturbation and virtual temperature correction(64, 59), the parcel buoyancy at a given height () is proportional to differences in sensible heat between the air parcel and the environment (), where is temperature perturbations of the air parcel (, with subscript “” referring to the air parcel) with respect to the environment () and J kg-1 K-1 is the the specific heat capacity of air at constant pressure. During an adiabatic process, it is common to assume that the air parcel conserves its moist static energy such that . Therefore, the vertical profile of or is given by
@@
For , the air parcel is saturated, such that the parcel specific humidity equals its saturated specific humidity (, with superscript “” referring to a quantity at saturation), which is a function of the parcel temperature and air pressure. As pressure difference between the air parcel and environment is negligible (64, 59), the difference in saturated specific humidity between the air parcel and environment is caused by difference in temperature, and thus can be written as a linear relation to the environmental saturated specific humidity () via . We further assume constant and approximate it by using Clausius-Clapeyron equation, where J kg-1 K-1 is gas constant for water vapor. Hence,
@@
Previous studies have explored the relationship between CAPE and environmental MSE, either under the radiative–convective equilibrium framework (129, 130, 54) or through linear regression methods (131). In contrast, our derivation begins with the fundamental definitions of parcel buoyancy and CAPE, similar to (132) but incorporating minimal assumptions about the geometric characteristics of CAPE, to establish a robust linear relationship between CAPE and key atmospheric parameters near the surface and in the free troposphere. The coefficient in Eq.12 may vary with , , and , which can be precisely adjusted using the full scaling equation (Eq.11) when applied to different regions or datasets. Here, the derived coefficient of 0.22 for high-CAPE environments over midlatitude land aligns closely with the linear regression results (0.22–0.29) reported for high CAPE over continental North America (131).
@@
The 3-hourly ERA5 reanalysis data (39, 40) from 1980–2022 are publicly available at https://rda.ucar.edu/datasets/ds633.0/.